# Wolfpack Adversarial Attack for Robust Multi-Agent Reinforcement Learning

### v2

@@ 5.2 Performance Comparison
Table 1 presents the average win rates over the last 100 episodes of MARL policies against attacker baselines. The results show that the proposed Wolfpack adversarial attack is significantly more powerful than existing methods like EGA and Random Attack. For example, EGA reduces Vanilla QMIX performance by $98.7-69.6=29.1\%$ and RANDOM by $99.3-76.9=22.4\%$ compared to the Natural scenario. In contrast, the Wolfpack attack reduces Vanilla QMIX performance by $98.7-39.4=59.3\%$ and RANDOM by $99.3-63.5=35.8\%$, demonstrating its greater impact. In addition, the proposed WALL framework, trained to defend against the Wolfpack attack, outperforms other robust MARL methods against all attack types, showcasing its superior robustness. Notably, despite RANDOM is trained against Random Attack and ROMANCE against EGA, WALL achieves better performance against both attack types. This result highlights the effectiveness of WALL in enabling robust learning under diverse attack scenarios. Fig. 6 further illustrates this in the 8m and MMM environments, where performance differences with existing methods are most pronounced, showing the average win rate of each policy over training steps under unseen Wolfpack adversarial attacks. The results reveal that WALL not only achieves higher robustness but also adapts more quickly to attacks. Similar trends are observed for other CTDE algorithms, such as VDN and QPLEX, as detailed in Appendix E.1, confirming the robustness of the proposed method.
@@ 5.4 Ablation Study
Figure 8: Component Evaluation of WALL and Wolfpack attack
@@
Figure 9: Number of Follow-up Agents ($m$)
@@
Figure 10: Step Selection Temperature ($T$)
@@ Appendix A Environmental Setup
All experiments in this paper are conducted in the SMAC (Samvelyan et al., 2019) environment, and this section provides a detailed description of its setup and features.
SMAC
@@
SMAC scenarios involve combat situations between an allied team controlled by learning agents and an enemy team managed by a scripted AI. These scenarios vary in complexity, unit composition, and terrain, challenging agents to use advanced micromanagement techniques such as focus fire, kiting, and terrain exploitation. Scenarios end when all units on one side are eliminated or when the time limit is reached. The objective is to maximize the win rate of the allied agents. Detailed descriptions of the scenarios and unit compositions are provided in Table A.1 and Figure A.1.
States and observations
In the SMAC environment, each agent receives partial observations, which contain information about visible allies and enemies within a fixed sight range of 9 units. These observations do not include global state information and are limited to the agent’s local view of the environment. Observations are specifically designed to support decentralized decision-making by individual agents, without access to global state features.
The global state, used during centralized training, aggregates the features of all agents and enemies. The state dimensionality depends on:
  • Ally state: A variable size based on the number of agents ($n_{\text{agents}}$), including their relative x (dim = $1$), relative y (dim = $1$), health (dim = $1$), energy (dim = $1$), shield (dim = $\mathrm{shield\_dim}$), and unit type (dim = $\mathrm{unit\_type\_dim}$). For each ally, the total feature size is $n_{\text{agents}}\times(4+\mathrm{shield\_dim}+\mathrm{unit\_type\_dim})$.
  • Enemy state: Similar to Ally state but excludes energy. For each enemy, the total feature size is $n_{\text{enemies}}\times(3+\mathrm{shield\_dim}+\mathrm{unit\_type\_dim})$.
  • Last action: Represents the most recent action taken by each agent. The total feature size is $n_{\text{agents}}\times n_{\text{actions}}$. Here, $n_{\text{actions}}$ refers to the number of actions.
These features are concatenated to form the observation vector for each agent. The dimensionality of the observation vector depends on the following factors:
  • Movement features: A fixed size representing available movement directions (dim = $4$).
  • Enemy features: A variable size based on the number of enemies ($n_{\text{enemies}}$), including their available action (dim = $1$), distance (dim = $1$), relative x (dim = $1$), relative y (dim = $1$), health (dim = $1$), shield (dim = $\mathrm{shield\_dim}$, where $\mathrm{shield\_dim}=1$ for Protoss units and 00 otherwise), and unit type (dim = $\mathrm{unit\_type\_dim}$, where $\mathrm{unit\_type\_dim}=0$ if there is 1 agent type, otherwise equals the number of agent types). For each enemy, the total feature size is $n_{\text{enemies}}\times(5+\mathrm{shield\_dim}+\mathrm{unit\_type\_dim})$.
  • Ally features: A variable size based on the number of allies ($n_{\text{allies}}$), where $n_{\text{allies}}=n_{\text{agents}}-1$, similar to Enemy features. For each ally, the total feature size is $n_{\text{allies}}\times(5+\mathrm{shield\_dim}+\mathrm{unit\_type\_dim})$.
  • Own features: Includes health (dim = $1$), shield (dim = $\mathrm{shield\_dim}$), and unit type (dim = $\mathrm{unit\_type\_dim}$). The total feature size is $1+\mathrm{shield\_dim}+\mathrm{unit\_type\_dim}$.
The precise dimensions of the observation and state vectors vary across different maps, as summarized in Table A.1.
@@
Map Ally Units Enemy Units State Dim Obs Dim Action Space
@@ 7 Marines 7 Marines
Table A.1: SMAC Scenarios with Units and Dimensions of State, Observation, and Action Space
(a) 3m
(b) 3s_vs_3z
(c) 2s3z
(d) 8m
(e) 1c3s5z
(f) MMM
Figure A.1: Visualization of SMAC Scenarios
@@ B.1 Practical Implementation of Planner Transformer
As shown in Fig. B.1, the estimated state $\hat{s}_{t+1}$ and observations $\hat{\mathbf{o}}_{t+1}$ are generated by the planning Transformer, which takes previous trajectories as input. Similarly, the estimated $Q$-difference $\hat{\Delta}Q_{t}^{\mathrm{WP}}$ is produced by the $Q$-diff Transformer, using previous states as input. The loss function $\mathcal{L}_{\mathrm{planning}}$ minimizes the prediction error for the next state $s{t+1}$ and observation $\mathbf{o}{t+1}$, while $\mathcal{L}_{\mathrm{qdiff}}$ minimizes the prediction error for $\Delta Q_{l}^{\mathrm{WP}}$ resulting from the Wolfpack attack. Figure B.1 illustrates the architectures of the transformers, where (a) represents the planning Transformer and (b) represents the $Q$-diff Transformer.
(a) Planning Transformer
(b) $Q$-difference Transformer
Figure B.1: Structure of Transformers
@@ B.2 Detailed Implementation of WALL
To achieve robustness, the training process minimizes the temporal difference (TD) loss, which incorporates the observed rewards, state transitions, and target $Q$-values parameterized by a separate target network, $\theta^{-}$. By leveraging this target network, the CTDE frameworks stabilize learning and mitigates the impact of the Wolfpack attack. The TD loss is defined as:
@@ C.1 Hyperparameter Setup
Adv Hypernet Embed Dim - - 64
Num Kernel - - 2
@@
Num Transformer Head 1
Num Transformer Decoder Layer 1
Transformer Embed Dim 64
@@
Num Total Attacks (Train) ($K$) 8 16 12 8 16 16
Num Total Attacks (Test) ($K$) 8 4 8 4 8 8
Num Wolfpack Attacks ($K_{\mathrm{WP}}$) 2 4 3 2 4 4
Num Follow-up Agents ($m$) 1 1 2 3 4 4
@@ E.1 Comparison Results for Other CTDE Algorithms
Method Map Name
2s3z
$K_{\mathrm{WP}}=2$
$t_{\mathrm{WP}}=3$
3m
$K_{\mathrm{WP}}=1$
$t_{\mathrm{WP}}=3$
3s_vs_3z
$K_{\mathrm{WP}}=2$
$t_{\mathrm{WP}}=3$
8m
$K_{\mathrm{WP}}=1$
$t_{\mathrm{WP}}=3$
MMM
$K_{\mathrm{WP}}=2$
$t_{\mathrm{WP}}=3$
1c3s5z
$K_{\mathrm{WP}}=2$
$t_{\mathrm{WP}}=3$
Mean
Natural Vanilla VDN $98.5\pm 1.35$ $96.0\pm 2.95$ $99.3\pm 0.50$ $97.8\pm 2.00$ $98.3\pm 0.50$ $83.5\pm 9.50$ $95.6\pm 1.65$
RANDOM $99.0\pm 0.10$ $\mathbf{99.8\pm 0.05}$ $99.4\pm 0.45$ $\mathbf{98.9\pm 0.05}$ $98.8\pm 1.00$ $97.6\pm 0.50$ $98.9\pm 0.36$
RARL $95.3\pm 0.50$ $92.5\pm 3.35$ $99.3\pm 0.45$ $96.3\pm 1.50$ $93.2\pm 1.50$ $91.8\pm 0.00$ $94.7\pm 0.55$
RAP $93.4\pm 1.30$ $96.8\pm 48.4$ $99.3\pm 0.45$ $98.7\pm 1.00$ $97.3\pm 0.50$ $97.3\pm 0.50$ $97.1\pm 0.37$
ERNIE $94.6\pm 3.85$ $99.4\pm 0.50$ $97.1\pm 0.47$ $98.4\pm 1.38$ $98.8\pm 0.82$ $96.5\pm 1.70$ $97.5\pm 0.86$
ROMANCE $98.2\pm 0.05$ $97.9\pm 1.05$ $99.4\pm 0.45$ $93.5\pm 5.35$ $97.9\pm 0.95$ $93.0\pm 3.10$ $96.6\pm 0.77$
WALL (ours) $\mathbf{99.8\pm 0.10}$ $99.4\pm 0.45$ $\mathbf{99.9\pm 0.10}$ $96.9\pm 2.00$ $\mathbf{99.4\pm 0.45}$ $\mathbf{100.0\pm 0.05}$ $\mathbf{99.2\pm 0.19}$
Random Attack Vanilla VDN $80.5\pm 1.50$ $65.5\pm 6.50$ $96.5\pm 0.50$ $52.5\pm 10.5$ $95.0\pm 1.00$ $76.5\pm 7.50$ $77.8\pm 1.92$
RANDOM $83.0\pm 1.00$ $\mathbf{94.5\pm 1.50}$ $96.0\pm 0.00$ $88.5\pm 0.50$ $95.5\pm 2.50$ $93.5\pm 1.50$ $91.8\pm 0.50$
RARL $81.0\pm 4.15$ $64.5\pm 8.40$ $97.3\pm 0.50$ $72.9\pm 0.05$ $76.3\pm 7.50$ $92.0\pm 0.20$ $80.7\pm 2.07$
RAP $89.8\pm 3.05$ $75.8\pm 37.9$ $97.7\pm 1.90$ $81.2\pm 0.50$ $95.2\pm 2.55$ $92.3\pm 0.50$ $88.7\pm 1.63$
ERNIE $80.2\pm 2.04$ $66.9\pm 2.84$ $93.1\pm 1.71$ $67.1\pm 4.92$ $89.1\pm 2.10$ $90.4\pm 4.54$ $81.1\pm 1.19$
ROMANCE $91.0\pm 5.00$ $79.0\pm 6.00$ $98.4\pm 0.40$ $54.0\pm 5.00$ $97.5\pm 0.50$ $92.0\pm 3.00$ $85.3\pm 0.62$
WALL (ours) $\mathbf{95.5\pm 1.50}$ $86.0\pm 6.00$ $\mathbf{99.4\pm 0.40}$ $\mathbf{90.0\pm 3.00}$ $\mathbf{98.5\pm 0.50}$ $\mathbf{98.1\pm 0.05}$ $\mathbf{94.6\pm 1.39}$
EGA Vanilla VDN $69.5\pm 7.50$ $54.5\pm 12.5$ $94.5\pm 1.50$ $59.5\pm 12.5$ $90.5\pm 2.50$ $74.5\pm 9.50$ $73.8\pm 1.00$
RANDOM $56.0\pm 6.00$ $73.0\pm 6.00$ $84.0\pm 13.0$ $84.5\pm 6.50$ $85.5\pm 1.50$ $86.5\pm 3.50$ $78.3\pm 2.42$
RARL $58.0\pm 1.15$ $79.0\pm 2.90$ $96.5\pm 0.70$ $76.7\pm 4.00$ $75.9\pm 8.20$ $81.3\pm 4.50$ $77.9\pm 1.86$
RAP $79.9\pm 3.05$ $93.0\pm 46.9$ $97.3\pm 0.45$ $85.7\pm 1.25$ $91.8\pm 3.75$ $87.2\pm 1.50$ $89.3\pm 0.21$
ERNIE $62.0\pm 14.9$ $62.5\pm 8.18$ $89.1\pm 6.58$ $74.7\pm 1.55$ $89.3\pm 2.51$ $82.2\pm 1.86$ $76.6\pm 3.62$
ROMANCE $86.5\pm 2.50$ $\mathbf{93.8\pm 3.00}$ $98.0\pm 0.00$ $76.5\pm 0.50$ $95.5\pm 2.50$ $92.0\pm 0.00$ $90.3\pm 0.58$
WALL (ours) $\mathbf{91.5\pm 0.50}$ $91.0\pm 2.00$ $\mathbf{99.0\pm 1.00}$ $\mathbf{90.0\pm 4.00}$ $\mathbf{97.5\pm 0.50}$ $\mathbf{95.5\pm 0.50}$ $\mathbf{94.1\pm 0.08}$
Wolfpack Adversarial Attack (ours) Vanilla VDN $54.0\pm 4.00$ $20.5\pm 5.50$ $91.5\pm 0.50$ $24.5\pm 12.5$ $18.5\pm 9.50$ $55.5\pm 2.50$ $44.1\pm 1.08$
RANDOM $47.0\pm 4.00$ $89.0\pm 1.00$ $90.0\pm 5.00$ $41.0\pm 14.0$ $18.5\pm 5.50$ $83.5\pm 0.50$ $61.5\pm 2.67$
RARL $59.5\pm 8.65$ $41.3\pm 16.4$ $96.8\pm 0.00$ $13.1\pm 2.00$ $24.3\pm 6.50$ $61.6\pm 11.7$ $49.4\pm 1.83$
RAP $64.3\pm 3.50$ $67.7\pm 33.8$ $98.9\pm 0.05$ $23.9\pm 6.05$ $53.3\pm 2.50$ $82.1\pm 5.40$ $65.0\pm 1.66$
ERNIE $33.1\pm 3.64$ $25.8\pm 6.48$ $93.0\pm 4.54$ $17.2\pm 9.46$ $23.5\pm 9.29$ $66.4\pm 13.5$ $43.2\pm 1.69$
ROMANCE $63.0\pm 10.0$ $46.5\pm 20.5$ $97.5\pm 0.50$ $19.5\pm 7.50$ $38.0\pm 5.00$ $83.0\pm 4.00$ $57.9\pm 3.08$
WALL (ours) $\mathbf{91.5\pm 2.50}$ $\mathbf{91.5\pm 2.50}$ $\mathbf{100.0\pm 0.00}$ $\mathbf{71.5\pm 1.50}$ $\mathbf{98.0\pm 1.00}$ $\mathbf{93.5\pm 4.50}$ $\mathbf{91.0\pm 0.83}$

### v3

@@ 5.2 Performance Comparison in MPE and SMAC
For SMAC, Table 2 presents the average win rates over the last 100 episodes of MARL policies under different attack baselines. The results show that the proposed Wolfpack adversarial attack is significantly more powerful than existing methods such as EGA and Random Attack. For example, EGA reduces the performance of Vanilla QMIX by $98.7-69.6=29.1\%$ and RANDOM by $99.3-76.9=22.4\%$ compared to the natural scenario. In contrast, the Wolfpack attack reduces Vanilla QMIX performance by $98.7-39.4=59.3\%$ and RANDOM by $99.3-63.5=35.8\%$, demonstrating its greater impact. In addition, the proposed WALL framework, which is trained to defend against the Wolfpack attack, outperforms other robust MARL methods under all attack types, showcasing its superior robustness. Notably, although RANDOM is trained specifically against Random Attack and ROMANCE against EGA, WALL achieves better performance against both attack types. These results highlight the effectiveness of WALL in enabling robust learning under diverse adversarial scenarios. Fig. 7 further illustrates this in the 8m and MMM environments, where performance differences with existing methods are most pronounced, showing the average win rate of each policy over training steps under unseen Wolfpack adversarial attacks. The results reveal that WALL not only achieves higher robustness but also adapts more quickly to attacks. Similar trends are observed for other CTDE algorithms, such as VDN and QPLEX, as detailed in Appendix E.1, confirming the robustness of the proposed method.
To support a more practical evaluation, we assess computational complexity and general robustness under common perturbations in Appendix E.4 and Appendix E.5, respectively. WALL incurs about 30% higher training cost than ROMANCE but achieves substantially better performance due to its critical step selection. For robustness, we consider perturbations including Gaussian noise in observations and test-time parameter shifts such as reduced allied HP, under which WALL still outperforms existing baselines. These results demonstrate the practical effectiveness of WALL under both computational and environmental challenges.
@@ 5.4 Ablation Study
Figure 9: Component evaluation of our WALL/Wolfpack attack
6 Limitations
While the proposed WALL significantly improves robustness in MARL, it has a few limitations. The first is the additional computational overhead introduced by training the Transformer for identifying critical steps. However, as shown in our analysis, this overhead is justified given that other baselines fail to achieve comparable performance even with extended training. Another limitation is the need for hyperparameter tuning to construct the Wolfpack attack. Nevertheless, the method is not highly sensitive to these hyperparameters, and the provided ablation study offers practical guidelines for selecting appropriate configurations.
@@ 6 Limitations
Figure 10: Number of follow-up agents $m$
@@
Figure 11: Step selection temperature $T$
@@ 7 Conclusions
Acknowledgment
This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2022-II220469, Development of Core Technologies for Task-oriented Reinforcement Learning for Commercialization of Autonomous Drones), Innovative Human Resource Development for Local Intellectualization program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2025-RS-2022-00156361), and Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2020-II201336, Artificial Intelligence graduate school support (UNIST)).
@@ Appendix A Environmental Setup
We conduct experiments in the MPE (Lowe et al., 2017) and SMAC (Samvelyan et al., 2019) environments. This section provides detailed descriptions of their setup and features.
A.1 Multi-Agent Particle Environments (MPE)
(a) PP_3/1
(b) PP_6/2
(c) PP_9/3
Figure A.1: Visualization of PP scenarios in the MPE environment
The Multi-Agent Particle Environment (MPE) (Lowe et al., 2017) is a widely used benchmark suite consisting of multi-agent scenarios. Agents are modeled as particles capable of movement and interaction, governed by simple physical dynamics. MPE includes both cooperative and competitive tasks, with each scenario sharing a continuous state space and typically partial observability. A standardized implementation of MPE is available through the PettingZoo library (Terry et al., 2021).
Scenarios
The MPE benchmark includes a variety of multi-agent scenarios. Among them, we focus on the predator-prey environment, which is well-suited for analyzing the impact of attacks on the cooperative structures among agents. We consider multiple variants denoted as $\texttt{PP}\_X/Y$, where $X$ represents the number of predator agents and $Y$ the number of prey agents. In all scenarios, prey agents follow a random policy. Detailed configurations for the three selected variants are summarized in Table A.1 and Fig. A.1.
State and observation spaces
Each agent in the MPE environment receives a partial observation, which includes its own position and velocity, the relative positions and velocities of other predators, and the relative positions of landmarks. The global state is constructed by concatenating the local observations of all agents.
Action space
The action space is discrete. Each agent can choose one of the four cardinal directions or do nothing.
Reward function
The reward function $R$ assigns a positive value to predator agents upon a successful collision with a prey:
$R$ $=\sum_{g\in\text{prey}}\mathbb{I}(\text{collision}(g,\text{predator}))$
where $\mathbb{I}(\cdot)$ is the indicator function, returning 1 if the condition inside is true and 0 otherwise. Here, $\text{collision}(i,j)$ denotes whether agent $i$ and agent $j$ are in physical contact.
Map Predators Prey State Dimension Obs Dimension Num. of Actions
PP_3/1 3 Agents 1 Agent 48 16 5
PP_6/2 6 Agents 2 Agents 156 26 5
PP_9/3 9 Agents 3 Agents 324 36 5
Table A.1: The number of agents, the dimensions of state and observation spaces, and the number of actions in MPE scenarios
A.2 The StarCraft Multi-Agent Challenge (SMAC)
(a) 3m
(b) 3s_vs_3z
(c) 2s3z
(d) 8m
(e) 1c3s5z
(f) MMM
Figure A.2: Visualization of SMAC scenarios
@@ A.2 The StarCraft Multi-Agent Challenge (SMAC)
SMAC scenarios involve combat situations between an allied team controlled by learning agents and an enemy team managed by a scripted AI. These scenarios vary in complexity, unit composition, and terrain, challenging agents to use advanced micromanagement techniques such as focus fire, kiting, and terrain exploitation. Scenarios end when all units on one side are eliminated or when the time limit is reached. The objective is to maximize the win rate of the allied agents. Detailed descriptions of the scenarios and unit compositions are provided in Table A.2 and Fig. A.2.
State and observation spaces
In the SMAC environment, each agent receives partial observations that contain information about visible allies and enemies within a fixed sight range of 9. These observations do not include any global state and are specifically designed to support decentralized decision-making based only on each agent’s local view of the environment.
The global state, which is used during centralized training, is constructed by aggregating the features of all agents and enemies. It consists of three primary components. The ally state includes each agent’s relative $x$ and $y$ positions, health, energy, shield (if applicable), and unit type. The enemy state is similar but excludes energy, containing relative positions, health, shield, and unit type of enemies. The last action component records the most recent action taken by each agent, represented as a one-hot encoded vector. The full global state is formed by concatenating these components, and its dimensionality depends on the number of agents, enemies, and available actions.
Each agent’s observation vector is separately constructed from the following elements. The movement features indicate the four cardinal directions the agent can move in, resulting in a fixed size of 4. The enemy features describe each observed enemy, including available action flag, distance to the agent, relative $x$ and $y$ positions, health, shield (if applicable), and unit type. The ally features encode the same types of information for all visible allies, excluding the observing agent. Finally, the own features contain the observing agent’s own health, shield, and unit type.
The precise dimensions of the observation and state vectors vary depending on the specific map scenario and unit composition, and are summarized in Table A.2.
@@
Map Ally Units Enemy Units State Dimension Obs Dimension Num. of Actions
@@ 7 Marines 7 Marines
Table A.2: The number of agents, the dimensions of state and observation spaces, and the number of actions in SMAC scenarios
@@ B.1 Practical Implementation of Planner Transformer
(a) Planning Transformer
(b) $Q$-difference Transformer
Figure B.1: Structure of Transformers
@@
As shown in Fig. B.1, the estimated state $\hat{s}_{t+1}$ and observations $\hat{\mathbf{o}}_{t+1}$ are generated by the planning Transformer, which takes previous trajectories as input. Similarly, the estimated $Q$-difference $\hat{\Delta}Q_{t}^{\mathrm{WP}}$ is produced by the $Q$-diff Transformer, using previous states as input. The loss function $\mathcal{L}_{\mathrm{planning}}$ minimizes the prediction error for the next state $s_{t+1}$ and observation $\mathbf{o}_{t+1}$, while $\mathcal{L}_{\mathrm{qdiff}}$ minimizes the prediction error for $\Delta Q_{l}^{\mathrm{WP}}$ resulting from the Wolfpack attack. Fig. B.1 illustrates the architectures of the Transformers, where (a) represents the planning Transformer and (b) represents the $Q$-diff Transformer.
@@ B.2 Detailed Implementation of WALL
To achieve robustness, the training process minimizes the temporal difference (TD) loss, which incorporates the observed rewards, state transitions, and target $Q$-values parameterized by a separate target network, $\theta^{-}$. By leveraging this target network, the CTDE frameworks stabilize learning and mitigate the impact of the Wolfpack attack. The TD loss is defined as:
@@ C.1 Hyperparameter Setup
Adv Hypernet Embed Dim. - - 64
Num. Kernel - - 2
@@
Num. Transformer Head 1
Num. Transformer Decoder Layer 1
Transformer Embed Dim. 64
@@
Scenario PP_3/1 PP_6/2 PP_9/3
Num. Total Attacks (Train) ($K$) 4 4 4
Num. Total Attacks (Test) ($K$) 4 4 4
Num. Wolfpack Attacks ($K_{\mathrm{WP}}$) 1 1 1
Num. Follow-up Agents ($m$) 1 3 5
@@
Num. Total Attacks (Train) ($K$) 8 16 12 8 16 16
Num. Total Attacks (Test) ($K$) 8 4 8 4 8 8
Num. Wolfpack Attacks ($K_{\mathrm{WP}}$) 2 4 3 2 4 4
Num. Follow-up Agents ($m$) 1 1 2 3 4 4
@@ Appendix E Additional Experiments Results
This section presents additional experimental results.  E.1 reports comparison results for other CTDE algorithms,  E.2 provides learning curves across additional SMAC scenarios,  E.3 presents a performance comparison for EGA,  E.4 discusses computational cost, and  E.5 includes general robustness experiments.
@@ E.1 Comparison Results for Other CTDE Algorithms
Method Scenario 2s3z 3m 3s_vs_3z 8m MMM 1c3s5z Mean
Natural Vanilla VDN $98.5\pm 1.4$ $96.0\pm 3.0$ $99.3\pm 0.5$ $97.8\pm 2.0$ $98.3\pm 0.5$ $83.5\pm 9.5$ $95.6\pm 1.7$
RANDOM $99.0\pm 0.1$ $\mathbf{99.8\pm 0.1}$ $99.4\pm 0.5$ $\mathbf{98.9\pm 0.1}$ $98.8\pm 1.0$ $97.6\pm 0.5$ $98.9\pm 0.4$
RARL $95.3\pm 0.5$ $92.5\pm 3.4$ $99.3\pm 0.5$ $96.3\pm 1.5$ $93.2\pm 1.5$ $91.8\pm 0.3$ $94.7\pm 0.6$
RAP $93.4\pm 1.3$ $96.8\pm 1.8$ $99.3\pm 0.5$ $98.7\pm 1.0$ $97.3\pm 0.5$ $97.3\pm 0.5$ $97.1\pm 0.4$
ERNIE $94.6\pm 3.9$ $99.4\pm 0.5$ $97.1\pm 0.5$ $98.4\pm 1.4$ $98.8\pm 0.8$ $96.5\pm 1.7$ $97.5\pm 0.9$
ROMANCE $98.2\pm 0.1$ $97.9\pm 1.1$ $99.4\pm 0.5$ $93.5\pm 5.4$ $97.9\pm 1.0$ $93.0\pm 3.1$ $96.6\pm 0.8$
WALL (ours) $\mathbf{99.8\pm 0.1}$ $99.4\pm 0.5$ $\mathbf{99.9\pm 0.1}$ $96.9\pm 2.0$ $\mathbf{99.4\pm 0.5}$ $\mathbf{100.0\pm 0.1}$ $\mathbf{99.2\pm 0.2}$
Random Attack Vanilla VDN $80.5\pm 1.5$ $65.5\pm 6.5$ $96.5\pm 0.5$ $52.5\pm 10.5$ $95.0\pm 1.0$ $76.5\pm 7.5$ $77.8\pm 1.9$
RANDOM $83.0\pm 1.0$ $\mathbf{94.5\pm 1.5}$ $96.0\pm 0.7$ $88.5\pm 0.5$ $95.5\pm 2.5$ $93.5\pm 1.5$ $91.8\pm 0.5$
RARL $81.0\pm 4.2$ $64.5\pm 8.4$ $97.3\pm 0.5$ $72.9\pm 0.1$ $76.3\pm 7.5$ $92.0\pm 0.2$ $80.7\pm 2.1$
RAP $89.8\pm 3.1$ $75.8\pm 37.9$ $97.7\pm 1.9$ $81.2\pm 0.5$ $95.2\pm 2.6$ $92.3\pm 0.5$ $88.7\pm 1.6$
ERNIE $80.2\pm 2.0$ $66.9\pm 2.8$ $93.1\pm 1.7$ $67.1\pm 4.9$ $89.1\pm 2.1$ $90.4\pm 4.5$ $81.1\pm 1.2$
ROMANCE $91.0\pm 5.0$ $79.0\pm 6.0$ $98.4\pm 0.4$ $54.0\pm 5.0$ $97.5\pm 0.5$ $92.0\pm 3.0$ $85.3\pm 0.6$
WALL (ours) $\mathbf{95.5\pm 1.5}$ $86.0\pm 6.0$ $\mathbf{99.4\pm 0.4}$ $\mathbf{90.0\pm 3.0}$ $\mathbf{98.5\pm 0.5}$ $\mathbf{98.1\pm 0.1}$ $\mathbf{94.6\pm 1.4}$
EGA Vanilla VDN $69.5\pm 7.5$ $54.5\pm 12.5$ $94.5\pm 1.5$ $59.5\pm 12.5$ $90.5\pm 2.5$ $74.5\pm 9.5$ $73.8\pm 1.0$
RANDOM $56.0\pm 6.0$ $73.0\pm 6.0$ $84.0\pm 13.0$ $84.5\pm 6.5$ $85.5\pm 1.5$ $86.5\pm 3.5$ $78.3\pm 2.4$
RARL $58.0\pm 1.2$ $79.0\pm 2.9$ $96.5\pm 0.7$ $76.7\pm 4.0$ $75.9\pm 8.2$ $81.3\pm 4.5$ $77.9\pm 1.9$
RAP $79.9\pm 3.1$ $93.0\pm 4.6$ $97.3\pm 0.5$ $85.7\pm 1.3$ $91.8\pm 3.8$ $87.2\pm 1.5$ $89.3\pm 0.2$
ERNIE $62.0\pm 14.9$ $62.5\pm 8.2$ $89.1\pm 6.6$ $74.7\pm 1.6$ $89.3\pm 2.5$ $82.2\pm 1.9$ $76.6\pm 3.6$
ROMANCE $86.5\pm 2.5$ $\mathbf{93.8\pm 3.0}$ $98.0\pm 0.2$ $76.5\pm 0.5$ $95.5\pm 2.5$ $92.0\pm 0.1$ $90.3\pm 0.6$
WALL (ours) $\mathbf{91.5\pm 0.5}$ $91.0\pm 2.0$ $\mathbf{99.0\pm 1.0}$ $\mathbf{90.0\pm 4.0}$ $\mathbf{97.5\pm 0.5}$ $\mathbf{95.5\pm 0.5}$ $\mathbf{94.1\pm 0.1}$
Wolfpack Adversarial Attack (ours) Vanilla VDN $54.0\pm 4.0$ $20.5\pm 5.5$ $91.5\pm 0.5$ $24.5\pm 12.5$ $18.5\pm 9.5$ $55.5\pm 2.5$ $44.1\pm 1.1$
RANDOM $47.0\pm 4.0$ $89.0\pm 1.0$ $90.0\pm 5.0$ $41.0\pm 14.0$ $18.5\pm 5.5$ $83.5\pm 0.5$ $61.5\pm 2.7$
RARL $59.5\pm 8.7$ $41.3\pm 16.4$ $96.8\pm 0.9$ $13.1\pm 2.0$ $24.3\pm 6.5$ $61.6\pm 11.7$ $49.4\pm 1.8$
RAP $64.3\pm 3.5$ $67.7\pm 33.8$ $98.9\pm 0.1$ $23.9\pm 6.1$ $53.3\pm 2.5$ $82.1\pm 5.4$ $65.0\pm 1.7$
ERNIE $33.1\pm 3.6$ $25.8\pm 6.5$ $93.0\pm 4.5$ $17.2\pm 9.5$ $23.5\pm 9.3$ $66.4\pm 13.5$ $43.2\pm 1.7$
ROMANCE $63.0\pm 10.0$ $46.5\pm 20.5$ $97.5\pm 0.5$ $19.5\pm 7.5$ $38.0\pm 5.0$ $83.0\pm 4.0$ $57.9\pm 3.1$
WALL (ours) $\mathbf{91.5\pm 2.5}$ $\mathbf{91.5\pm 2.5}$ $\mathbf{100.0\pm 0.0}$ $\mathbf{71.5\pm 1.5}$ $\mathbf{98.0\pm 1.0}$ $\mathbf{93.5\pm 4.5}$ $\mathbf{91.0\pm 0.8}$